# One-Shot Real-World Demonstration Synthesis for Scalable Bimanual Manipulation

### v1

@@ D.3 Two More Complex, Long-Horizon Tasks
D.4 Few-Shot-Based Demonstration Synthesis
While BiDemoSyn is designed around a one-shot paradigm, the framework naturally extends to a few-shot setting, where multiple real demonstrations may serve as complementary seeds. This subsection studies the effect of incorporating a small number of additional demonstrations (5–20 samples) obtained through standard teleoperation or kinesthetic teaching (used in this study).
Few-Shot Integration into BiDemoSyn. Given multiple seed demonstrations, we incorporate them in two ways: (1) Enhanced Object Variation Coverage. Additional demonstrations expand the diversity of observed object geometries, grasp points, and intermediate contact poses. These are incorporated into the variant blocks during trajectory adaptation. (2) Improved Robustness of Reference Motion. Multiple demonstrations allow BiDemoSyn to aggregate or select more reliable canonical motion patterns for the invariant blocks, mitigating single-demonstration idiosyncrasies. Under these two settings, the followed overall synthesis pipeline remains unchanged.
Experimental Findings. We have summarized results of this ablation study in Tab. 2. These new experiments show: (1) In-distribution (ID) performance. Adding few-shot data provides little to no improvement. Policies trained with BiDemoSyn’s one-shot–generated datasets already exhibit near-saturation performance for seen objects and placements. (2) Out-of-distribution (OOD) performance. Few-shot data provides moderate gains, primarily by increasing the number of observed object instances and geometric variations, thereby improving generalization to unseen shapes and dimensions. These findings align with the intuition that BiDemoSyn already produces high-quality, physically faithful demonstrations from a single demonstration (refer Fig. 5). Additional few-shot demonstrations mainly help by expanding the representational coverage rather than correcting failure cases caused by hardware limitations of the embodied robot entity itself.
Additional Conclusion. The few-shot extension confirms that BiDemoSyn supports scalable integration of multiple demonstrations, maintains strong performance in the one-shot setting, benefits from few-shot seeds mostly in the OOD regime, and requires no modification to its core synthesis pipeline. This suggests a promising avenue for future research: exploring hybrid one-shot + few-shot paradigms that balance data-collection cost with maximal generalization capability.
D.5 VFM-based Object Anchoring in Clutter
To evaluate the applicability of BiDemoSyn under more realistic and visually complex environments, we additionally examine its object-anchoring capability when the target objects are placed in cluttered scenes with distractors, varied background patterns, and partial occlusions. Although the main experiments intentionally isolate environmental variables to study the core challenge of real-world demonstration synthesis, BiDemoSyn is fully compatible with more complex perception conditions due to its reliance on Vision Foundation Models (VFMs) for object segmentation.
Clutter-Aware Object Localization. We employ Florence-2 xiao2024florence and SAM2 ravi2024sam to detect and segment the task-relevant items (e.g., bottle, mug) using either text queries or exemplar-based prompts. In cluttered tabletop scenes, the VFM stack consistently produces high-quality masks, even when distractors share similar colors or shapes. Qualitatively, the predicted masks remain stable across moderate occlusions, enabling the system to isolate the target object with minimal ambiguity. Fig. 17 illustrates representative examples where both bottle and mug are correctly segmented despite surrounding tools, utensils, packaging, and nonrigid paper balls.
Implications for Demonstration Synthesis. Because BiDemoSyn’s variant-block adaptation depends only on (1) reliable 2D masks and (2) centroid- and principal-axis–based orientation inference, accurate object anchoring is the only requirement for extending the system to cluttered environments. Our experiments show that the VFM-based perception pipeline provides sufficiently robust segmentation to support (a) pose-aligned trajectory modulation and (b) collision-aware grasp adjustments, even when the background is visually nonuniform.
Limitations and Future Directions. While VFMs significantly enhance robustness to visual complexity, extreme occlusions or heavy inter-object entanglement may still degrade performance. Extending BiDemoSyn to handle these cases will require multi-view sensing, active view planning, or integrating VLM-driven object-query disambiguation. Nevertheless, these initial results demonstrate that VFM-based anchoring provides a practical and reliable foundation for deploying BiDemoSyn beyond clean tabletop settings, thereby enabling future extensions to household, warehouse, and mobile manipulation scenarios.

### v2

@@ V-B Comparison and Presentation of Results
($\mathsf{A}3$): Policies trained on BiDemoSyn data can achieve better generalization to unseen variations. To evaluate the generalization to novel instances, we conduct controlled tests across all six tasks. For each task, we exclude demonstrations associated with one or a pair of randomly selected objects (e.g., an arbitrary bottle in unscrew or a bottle-cup pair in pouring) from the training set, resulting in training sizes of 2916/3888/1008/972/324/756 demos for six tasks, respectively. These excluded objects are reintroduced exclusively during real-world testing, with identical data splits applied to DemoGen and YOTO baselines for fair comparison. As shown in Tab. I right, policies trained on BiDemoSyn data consistently achieve higher success rates compared to all baselines in out-of-distribution (OOD) settings. While all methods exhibit performance drops relative to ID evaluations, BiDemoSyn’s superior OOD robustness highlights its ability to capture task-invariant features through vision-aligned synthesis. To further analyze spatial generalization, we vary the density of positional and orientation coverage in training data: for unscrew, we synthesize trajectories with sparse-to-dense workspace coverage (see Fig. 5A), and for reorient, we incrementally increase the diversity of object orientations (see Fig. 5B). To exclude the effect of data size differences, the total number of train-set is always augmented to a standard size (e.g., #1152 and #1008) via the DemoGen to edit some observations at small distances. Finally, policies trained on BiDemoSyn data with dense spatial coverage (e.g., $10\times 10$ workspace coverage in unscrew) achieve 76.7% success on unseen positions, compared to 53.3% for sparse coverage ($6\times 6$ grid cells). Similarly, orientation diversity in reorient improves generalization from 37.3% (lower diversity with #Ornt.=3) to 73.3% (higher diversity with #Ornt.=7). These results confirm that BiDemoSyn’s ability to cheaply synthesize positionally and orientationally diverse demos directly translates to enhanced policy generalization, aligning with empirical insights that broad data coverage mitigates spatial distribution shifts.
Qualitative Results: Fig. 6 shows effectiveness across six tasks. For example, the plugpen policy precisely aligns and inserts the pen into the narrow cap even with $\pm$5mm positional noise. In inserting, dual-arm coordination achieves rough but proximate peg-in-hole precision. The unscrew demonstrates robust rotational synchronization, successfully loosening lids with variable thread tightness. The pouring policy adapts liquid flow control to container geometries, minimizing spills despite tilted orientations. For pressing, predicted actions reliably activate nozzles without over-pressuring, while reorient handles object handovers and flips with smooth dual-arm transitions. These results corroborate our quantitative findings, emphasizing BiDemoSyn’s ability to handle real-world complexity in contact-rich cases.
Figure 7: Visualization for the cross-embodiment deployment of BiDemoSyn-trained policies on an unseen humanoid-style dual-arm robot for the plugpen and reorient tasks.
($\mathsf{A}4$): BiDemoSyn can seamlessly integrate with few-shot demonstrations to improve robustness. Although BiDemoSyn is formulated in a one-shot paradigm, it can seamlessly incorporate additional real demonstrations (5–20 samples) as complementary seeds without altering the synthesis pipeline. We summarized this ablation results in Tab. II. Empirically, we can find that few-shot extensions provide limited gains under ID settings, where one-shot–generated datasets already yield near-saturated performance on seen settings. In contrast, moderate but consistent improvements are observed under OOD conditions, primarily due to expanded coverage of object geometries and intermediate contact states introduced by new demonstrations. These results suggest that BiDemoSyn’s one-shot synthesis already produces high-quality, physically grounded demonstrations, while few-shot seeds mainly contribute by enriching representational diversity rather than correcting systematic failure modes. Overall, this study confirms that BiDemoSyn offers a favorable trade-off: it remains highly effective in the one-shot regime, benefits selectively from few-shot data in OOD scenarios, and supports hybrid one-shot + few-shot usage without increasing complexity.
($\mathsf{A}5$): Policies trained on BiDemoSyn data transfer effectively across robot embodiments in a zero-shot manner. BiDemoSyn facilitates zero-shot cross-embodiment policy deployment by design, owing to its object-centric observation formulation and the use of end-effector–centric 6-DoF action representation. These choices decouple policy learning from robot-specific kinematics and enable direct transfer to new hardware platforms. To evaluate this capability, we deploy visuomotor policies trained on the primary dual-arm platform to a previously unseen humanoid-style bimanual robot, without additional training or fine-tuning. We consider two representative tasks, plugpen and reorient, using EquiBot policies trained under ID settings. As shown in Fig. 7, the transferred policies successfully reproduce precise bimanual alignment for insertion and reorientation. Quantitative evaluations indicate that success rates on the new embodiment are comparable to those reported in Tab. I left. These results demonstrate that BiDemoSyn-trained policies exhibit strong embodiment-agnostic behavior, highlighting the potential for scalable deployment across heterogeneous robotic platforms.
VI Conclusion and Limitation
We introduce BiDemoSyn, a real-world demonstration synthesis framework that enables scalable bimanual manipulation learning from a single kinesthetic example. By decomposing demonstrations into invariant coordination blocks and object-dependent adaptations, and instantiating them via vision-based object anchoring and trajectory optimization, BiDemoSyn generates large-scale, physically feasible demonstrations without simulation or repeated teleoperation. Across diverse bimanual tasks, policies trained on BiDemoSyn data achieve strong generalization to novel objects, long-horizon compositions, and unseen robot embodiments. We further show that the framework naturally extends to few-shot synthesis, improving robustness with minimal additional supervision. This work establishes a new paradigm for scalable imitation learning, bridging the divide between data quantity and physical fidelity in bimanual robotic manipulation.
Limitations: While BiDemoSyn excels in static, geometrically varied scenes, it faces challenges in fully dynamic environments and extreme shape variations (e.g., highly deformable or articulated objects). The current hierarchical optimization assumes quasi-static motions, limiting its applicability to highly dynamic tasks like catching. Additionally, synthesizing trajectories for multi-stage tasks with interdependent contacts (e.g., lacing ropes) requires further extension. Future work will integrate dynamic perception and adaptive contact modeling to address these constraints.
Appendix
This supplementary provides detailed explanations and extensions to the main paper. Sec. VII outlines the design rationale of six bimanual manipulation tasks, including platform specifications, task-specific objectives, and one-shot demonstration acquisition protocols. Sec. VIII details the data collection standards, UI-assisted collection tools, and post-processing workflows. These constitute the trajectory synthesis pipelines for generating real-world demonstrations with BiDemoSyn. We have also added more details about BiDemoSyn to help to understand it. Sec. IX elaborates on policy training, covering implementation specifics for baselines (DemoGen [93], YOTO [107]) and diffusion policies (DP3 [98], EquiBot [94]), alongside real-robot deployment. Sec. X provides additional visualizations of real-robot executions and failure case analyses, offering insights into current limitations. Sec. XI concludes with reflections and future directions, aiming to extend BiDemoSyn to challenge more complex tasks and enhance cross-domain generalization.
VII Design of Bimanual Manipulation Tasks
VII-A Hardware and Platform
[…]
Figure 8: (Top) Primary. The fixed-base dual-arm platform having a table with two robot arms, two grippers and the binocular camera. (Bottom) Auxiliary. The humanoid dual-arm platform for validating cross-embodiment capabilities.
VII-B Bimanual Task Formulation
We design six bimanual tasks (plugpen, inserting, unscrew, pouring, pressing and reorient) to comprehensively evaluate dual-arm coordination and generalization. Each task involves at least two category-level object instances (Fig. 9) and requires diverse primitive skills, such as single-arm actions (grasping, placing, precision rotation) and dual-arm coordination (fixed-twisting, plug-in, handover). All tasks are inherently long-horizon, starting from goal-conditioned initial grasping (objects fully separated from robots), contrasting prior works already grasping/holding the manipulated object and focusing on short-horizon atomic skills such as untwisting [54, 1] or handover [37]. Fig. 10 shows some examples. Below we detail each task:
Figure 9: Object assets involved in six bimanual manipulation tasks. All objects have been scaled down proportionally.
  • pressing: press bottle and catch water using the cup. States: A shampoo bottle (with a pressable nozzle) containing water and an upright empty handle-free cup are placed apart. Steps: Right arm vertically presses the nozzle; left arm tilts the cup to catch water. Challenges: Nozzle positioning accuracy ($\pm$5mm), open-loop fine-grained force control.
[…]
VII-C Obtaining and Decomposing the One-Shot Demonstration
To enable reliable downstream synthesis, the one-shot kinesthetic demonstration must be collected in a way that naturally exposes task-relevant structure. During kinesthetic teaching, the demonstrator physically guides the two manipulators through the task while pausing at intuitive semantic boundaries, such as after a stable grasp, after completing an alignment adjustment, or before switching from transportation to interaction. These pauses, which arise organically from human motion patterns, produce clear temporal discontinuities that align well with the block-based representation introduced in the main paper. Importantly, this process does not require any manual annotation or labeling. The demonstration is captured in a continuous stream exactly as in standard kinesthetic teleoperation, and the boundaries between coarse blocks emerge directly from the dynamics of human-guided motion. This makes the procedure no more labor-intensive than collecting a single ordinary demonstration.
We explored alternative strategies for fully automatic segmentation, such as keyframe/keypose detection and curvature-based change-point identification [39, 78, 61]. However, reliably identifying semantically meaningful waypoints from a single demonstration proved brittle, especially when waypoints must correspond to subtle but crucial phases of contact establishment, re-grasping, or dual-arm synchronization. Hardware-assisted approaches (e.g., instrumented gloves as in DexCap [82] or end-effector teaching interfaces as in UMI [17] and DexUMI [92]) can offer additional cues for segmentation but introduce alignment overhead, cross-morphology calibration challenges, and potential loss of semantic correspondence between human motion and robot embodiment. Given these trade-offs, the kinesthetic-teaching-driven segmentation remains a pragmatic and robust choice: it imposes no additional annotation burden, preserves semantic alignment across blocks, and provides boundaries well suited for constructing invariant and variant components for real-world synthesis.
VIII Details of Data Collection and Process
Figure 11: The grid cell division for each task. For tasks involving two manipulated objects, the total number of grid cells will be divided equally between the left and right side.
VIII-A Task-Oriented Collection Standards
[…]
Figure 12: UI interfaces. When collecting data, six tasks need to display different grid cells in real time, as well as perceive and track task-related objects. Note that these points, crosses and lines are drawn digitally, which are not marks in the real world.
VIII-B Collection UI Interface Development
To streamline data collection, we develop a real-time UI interface (Fig. 12) that visualizes task-specific grid cells, dynamically tracks object positions/poses, and validates alignment with the one-shot demonstration. The interface overlays grid layouts onto live camera feeds, highlights detected objects (using YOLOE [81] for fast detection and segmentation), and marks completed cells using red crosses. The position of objects is represented by red dots (some objects also use yellow arrows to indicate their orientation). In addition, for cups and bottles that do not need to pay attention to changes in orientation, we did not use the mask centroid as their representive point during collection, but instead used the lowest endpoint in contact with the desktop. A sidebar displays positional/orientational deviations from the reference demonstration, enabling quick verification. If object perception fails, the system falls back to Florence2 [89] and SAM2 [74] for robust detection and segmentation. Operators trigger a capture-align-validate-save cycle every 2 seconds. And successful saves prompt visual feedback (UI flash of corresponding grid cells), signaling readiness for the next sample. This setup eliminates manual grid marking and minimizes human idle time, achieving a high speedup while ensuring spatial consistency. Please refer our Supplemental Videos that demonstrate the vivid workflow, illustrating how the UI balances efficiency with precision.
VIII-C Post-Processing of Collected Data
This part includes the last two stages of BiDemoSyn: initial frame alignment and trajectory optimization. First, pre-recorded positional/orientational deviations (validated via the UI interface during collection) align novel observations with the one-shot demonstration reference frame using the rigid transformation. Next, trajectory optimization ensures kinematic feasibility. We prune singular configurations and validate collision-free paths using the robot IK solvers and motion planners. Over 96% of trajectories pass validation, with failures (e.g., unreachable corner grid cells) either re-collected or discarded. Finally, validated trajectories are stored in $\mathcal{D}_{syn}$ to complete the synthesis pipeline.
VIII-D Hyperparameters and Predefined Symbols
In the method section, several thresholds and parameters are introduced for clarity and reproducibility. Here we further explain their meaning and selection strategy. For instance, threshold symbols $\delta$, $\zeta$ and $\gamma$ in Eqn. 4, Eqn. 5 and Eqn. 10 are predefined only for one-shot deconstruction. The symbol $\delta$ is a minimal motion saliency threshold, used to identify boundaries between coarse blocks. The symbol $\zeta$ is a static tolerance threshold to segment motion boundaries under quasi-static assumptions. The symbol $\gamma$ is a refinement threshold used to create fine-grained AEPs after bounding block duration ($T_{\textrm{\scriptsize max}}$). And the scaling factor $\lambda$ in Eqn. 10 controls the adaptation from the canonical one-shot demonstration to new object instances with varying sizes. Theoretically, $\lambda\in\mathbb{R}^{+}$, yet empirical tuning shows that values within $[0.8,1.0]$ are sufficient for stable grasp transfer without destabilizing contact. Larger or smaller values tend to shift grasp points away from functional regions (e.g., bottle mid-section in unscrew), while $\lambda=1.0$ already suffices for most tasks. Similarly, thresholds for collision-checking or pose-alignment rejection (e.g., rejecting infeasible IK solutions or near-singular samples) were determined through small-scale grid search and validated in pilot runs. Importantly, such hyperparameters are not per-trial tunable knobs, but rather fixed system constants, ensuring consistent data synthesis across tasks.
VIII-E Evaluation of Intermediate Processes
We provide further evaluation to justify intermediate design choices from three aspects: Segmentation Quality, Pose Estimation Robustness and Impact on Data Reliability. (1) Firstly, object masks were obtained from Florence2 [89] and SAM2 [74], two state-of-the-art vision foundation models (VFMs). On held-out captures without foreground occlusion, the segmentation accuracy exceeded 97%, ensuring reliable geometric inputs for subsequent processing. (2) Secondly, the axis alignment for objects with a distinct orientation was based on classical image-moment theory: first-order moments define centroids, second-order moments define orientation. This method, given reliable masks, yields highly stable results. Quantitative inspection of $\sim$200 sampled trials revealed <<2% pose estimation errors, and only under severe occlusion. (3) Thirdly, in Sec. X-D, we showed that perception and orientation errors together account for $\sim$47% of failures, highlighting their importance. Nevertheless, these intermediate modules are already close to optimal under current open-loop constraints, which explains the high success rate ($\sim$98%) of the data collection system.
VIII-F Positioning with Respect to Prior Works
To avoid potential confusion with prior related works such as ReKep [38], ODIL [84] and MAGIC [60], we emphasize both the similarities and the fundamental differences. Like these methods, BiDemoSyn exploits one-shot demonstrations and compositional reasoning; however, its core objective is not direct policy execution, but rather rapid and scalable synthesis of real-world training demonstrations. This distinction is crucial: while prior methods target zero-/few-shot policy generalization (often trading reliability for flexibility), BiDemoSyn deliberately focuses on maximizing data quality and efficiency, enabling downstream training of robust visuomotor policies (e.g., DP3 [98] and EquiBot [94]). Our evaluation against DemoGen [93] and YOTO [107], which are two strong baselines addressing the same bottleneck, demonstrates that BiDemoSyn achieves a superior balance between efficiency and data reliability, which is the practical need for scaling imitation learning.
A concurrent work MoMaGen [49] tackles scalable bimanual mobile manipulation by generating diverse trajectories in simulation, focusing on base reachability and camera visibility through constrained optimization. Although sharing a high-level goal of reducing human effort, MoMaGen and BiDemoSyn operate under fundamentally different assumptions: MoMaGen relies on noise-free simulation assets and virtual scene perturbations, whereas BiDemoSyn performs real-world demonstration synthesis, directly addressing physical-domain challenges such as noisy perception, contact feasibility, and one-shot decomposition under real sensor observations. As a result, MoMaGen expands diversity through simulated scene variation, while BiDemoSyn enables physically grounded generalization across spatial and category-level object variations without sim-to-real transfer. Thus, BiDemoSyn should be viewed not as a direct competitor to simulation-based frameworks like MoMaGen, but as a complementary real-world solution specifically designed for scalable, high-fidelity data generation where simulation pipelines cannot fully capture real-world physical constraints.
VIII-G Evaluation of Demonstration Quality
Visual Fidelity. This dimension evaluates whether the observations included in a synthesized demonstration faithfully reflect the real physical scene without introducing artifacts. Teleoperation and BiDemoSyn preserve raw RGB-D observations, ensuring perfect consistency with the physical environment. YOTO [107] and DemoGen [93] modify the seed point cloud to reposition objects, which introduces geometric distortions and depth inconsistencies, particularly as the displacement grows (see Fig. 4B in the main paper). These artifacts degrade the policy’s perception module by coupling unrealistic geometry with real sensor noise.
Trajectory Feasibility & Physical Grounding. A demonstration is considered high-quality if the associated action sequence: (1) yields a collision-free, dynamically stable trajectory, (2) respects the contact geometry implied by the task semantics, and (3) can be executed on the real robot without additional correction or replanning. Teleoperation and YOTO [107] provide ground-truth trajectories. BiDemoSyn produces partially reused (invariant blocks) and partially adapted (variant blocks) trajectories aligned with real-world object pose and geometry. DemoGen [93] synthesizes trajectories directly from edited scene representations, which can deviate from physically stable contact geometries.
Finally, Fig. 4A in the main paper reflects the interplay of these two components. Demonstrations with high-quality observations and physically feasible actions rank highest. This ordering also predicts downstream policy performance: policies trained with cleaner, physically grounded demonstrations demonstrate higher success rates and better robustness under OOD evaluation. These expanded analyses clarify the qualitative and quantitative factors driving the reported comparisons and support the role of demonstration quality as a key determinant of real-world imitation learning performance.
VIII-H Diversity of Synthetic Demonstrations
A natural question is whether additional diversity (such as reordering or randomizing block sequences) would improve the effectiveness of the synthesized dataset. In BiDemoSyn, we intentionally do not introduce such sequence-level variations. Our empirical findings align with recent analyses in large-scale teleoperation datasets focusing on the diversity exploration and ablation for scalable robotic manipulation [77], which show that low-level behavioral variability (e.g., minor differences in grasp location, end-effector speed profiles, or redundant micro-motions) does not necessarily benefit policy learning and may even hinder convergence. Instead, we adopt the principle that task success is primarily determined by achieving correct interactions, not by how those interactions are executed.
Accordingly, in BiDemoSyn’s design, we maintain the original causal structure of the task by preserving the block ordering extracted from the one-shot demonstration, while introducing diversity along the dimensions that matter for generalization, namely: (1) spatial variation of object pose (translation + rotation), and (2) category-level instance variation (shape and size differences within the same object class). These forms of variation are directly tied to visuomotor generalization and produce clear benefits in downstream policy performance. In contrast, randomizing the internal block structure introduces behavioral diversity that is orthogonal to task semantics and does not improve success rates in our preliminary tests. We therefore avoid unnecessary variation and instead focus on environment-induced diversity, which is both meaningful and consistent with the underlying task geometry.
Figure 13: The adapted visuomotor network architecture for imitation learning of bimanual tasks. The input visual observation is simplified to a point cloud of manipulated objects. Note that the vision encoder is slightly different in DP3 [98] and EquiBot [94]. The illustrated one is for EquiBot.
Figure 14: Qualitative real robot rollout samples for defined six bimanual tasks. Best to view after zooming in.
Figure 15: Qualitative real robot rollout examples for two newly added long-horizon tasks (pouring and reorient+unscrew).
Figure 16: Qualitative examples of applying dynamic interferences during the pre-grasping execution. From top to bottom, they are segments of the dynamic closed-loop grasping phase of tasks pouring and reorient+unscrew, where each object is manually disturbed from one to three times. The red arrow indicates the direction of the manually moved object (interfering). The cyan arrow and yellow arrow indicate the movement direction of the left and right robotic arms (chasing), respectively.
IX Training and Deployment of Imitation Policies
IX-A Reproduction of Baselines
[…]
Figure 17: Qualitative real robot rollout samples for two tasks (plugpen and reorient) on the auxiliary dual-arm platform.
IX-B Implementation of Imitation Policies
Both DP3 [98] and EquiBot [94] (refer Fig. 13) ingest segmented 3D object point clouds but differ in network architectures. Specifically, DP3 employs specially designed multiple MLPs to encode point clouds into latent features, then directly regressing dual-arm keyposes. EquiBot enhances robustness via a SIM(3)-equivariant PointNet variant: point clouds undergo rotation/scale-invariant feature extraction before action prediction. This equips EquiBot with improved generalization to unseen object variations, achieving 5–12% higher success rates than DP3 under identical training data and settings. Both policies operate in open-loop, aligning with synthesized demonstration formats. Below, we give a more detailed description of the visuomotor policy modification for the dual-arm action prediction.
Network Architecture. For DP3, it is a variant of Diffusion Policy [16] with a simpler point cloud encoder. It also designs a two-layer MLP to encode robot proprioceptive states before concatenating with the observation representation. For EquiBot, we use a SIM(3)-equivariant PointNet++ [95] with 4 layers and hidden dimensionality 128 as the feature encoder. For the noise prediction network, we inherits hyperparameters from the original Diffusion Policy. Specifically, to optimize for inference speed in all experiments, we use the DDIM scheduler [79] with 8 denoising steps, instead of the DDPM scheduler [35] which performs up to 100 denoising steps.
In addition, we have adopted the object-centric point cloud input. At inference time, we also need to preprocess the binocular RGB observations to obtain the point cloud of manipulated objects. This core design relies on the still rapidly developing capabilities of vision foundation models (VFMs). Here we leverage SOTA open vocabulary detection method Florence-2 [89] and segmentation method SAM2 [74] to automatically extract object masks and then filter out corresponding point clouds. Despite this, occasionally we may fail to segment desired objects accurately, and in these special cases we will manually correct the masks. These cases are not counted as failed evaluation trails due to not involving significant elements of bimanual robot manipulation. Because we believe that the next generation of VFMs can alleviate these problems, or we can directly address them through domain adaptation, test-time adaptation, or adjusting input prompts.
X More Results Visualization and Analysis
X-A More Examples from Real Rollouts
Fig. 14 supplements more qualitative results with additional real-robot execution visualizations, highlighting nuanced state transitions and task-specific challenges. These visualizations underscore BiDemoSyn’s ability to handle real-world complexities (such as mechanical tolerances and imperfect perception), while also exposing limitations in dynamic force modulation (e.g., over-pressing bottles or lids). Please refer our Supplemental Videos which provide frame-by-frame analyses of these executions, further dissecting critical phase and offering insights into policy decision-making under uncertainty.
X-B Two More Complex Long-Horizon Tasks
To further evaluate the scalability of BiDemoSyn beyond the introduced six bimanual tasks, we add two new long-horizon tasks that require sequential manipulation steps and accumulate more physical error throughout the execution:
To utilize BiDemoSyn, these two newly added tasks were synthesized following the same three-stage pipeline described in the main text. We briefly summarize the configurations: for reorient+unscrew, we collected initial observations for 4 bottle instances, each placed at 7 orientations across 36 grid cells, yielding 1008 synthesized demonstrations; for unscrew+pouring, we collected initial observations for 2 bottle and 2 cup instances, each placed on 18 grid cells, yielding 1296 synthesized demonstrations. Importantly, both tasks introduce longer temporal horizons, sequential coordination, and increased compounding error, thus providing a stronger stress test than the original six tasks.
Then, using the synthesized datasets, we trained visuomotor policies for each long-horizon task based on the adapted EquiBot architecture. Across both tasks, the trained policies exhibited high robustness to spatial variation (different placements and bottle orientations), reliable execution of multi-step dual-arm coordination, and consistent task success under mild external perturbations introduced during pre-grasp interpolation. In Fig. 15, we present some qualitative results with additional real-robot execution visualizations of these two new tasks. And in Fig. 16, we provide illustrative examples of continuous dynamic interference scenarios among tasks pouring and reorient+unscrew. Full dynamic demonstrations and rollouts of all eight tasks (previous 6 tasks and the new 2 tasks) are included in the Supplemental Videos. These results further validate BiDemoSyn’s ability to rapidly scale to new, long-horizon tasks with minimal additional engineering, and reinforce its suitability as a practical framework for synthesizing large-scale, real-world training data.
Figure 18: Qualitative examples of applying dynamic interferences in pre-grasping for the task reorient on the auxiliary dual-arm platform, where the object is manually disturbed up to five times. The red arrow indicates the direction of the manually moved object (interfering). The yellow arrow indicates the movement direction of the right robotic arm (chasing).
X-C Cross-Embodiment Transferability Testing
Beyond training and evaluating policies on the primary dual-arm platform (Fig. 8 top), we further investigate the cross-embodiment transferability of policies trained with BiDemoSyn. Specifically, we deploy the learned visuomotor policies on a distinct humanoid-style dual-arm robot with different kinematic structures, arm mounting configurations, and workspace geometry (Fig. 8 bottom). Importantly, no additional demonstrations, fine-tuning, or embodiment-specific retraining are introduced.
We evaluate two representative bimanual tasks—plugpen and reorient—which respectively emphasize precise dual-arm alignment and large-range end-effector rotation. As illustrated in Fig. 17, policies trained on the original platform can be directly executed on the new robot after a simple coordinate transformation of the predicted 6-DoF end-effector poses, without requiring explicit calibration of dynamic or contact parameters. Across extensive trials, both tasks achieve stable and consistent execution performance, comparable to those reported on the source platform. To further assess robustness, we introduce external perturbations during the pre-grasp stage (similar to Fig. 16), including manual displacement and rotation of target objects. As shown in Fig. 18, the policy can reliably recover from such disturbances and successfully complete the task, demonstrating that the transferred policy preserves not only task semantics but also robustness to moderate environmental variations. These qualitative results highlight a key advantage of BiDemoSyn: by adopting object-centric observations and embodiment-agnostic action representations, the resulting policies exhibit strong zero-shot transferability across heterogeneous robot embodiments, suggesting promising scalability to broader robotic platforms. More rollouts on this humanoid platform are in our Supplemental Videos.
Figure 19: (Top) Failure cases in the task plugpen that wrongly detected the marker body, and the task reorient+unscrew that incorrectly estimated the lying bottle’s orientation. (Bottom) Failure cases in the task inserting where the robotic arm’s end effector occasionally obscures the marker pen, and the task reorient where human interference causes the hand to partially obscure the manipulated spoon/shovel.